A dealer wishes to produce 300 gallons of oil worth 40 cents a quart. He has oil worth 36 cents a quart and oil worth 52 cents a quart. How many quarts of each should he mix?
Answer by lwsshak3(11628)

A dealer wishes to produce 300 gallons of oil worth 40 cents a quart. He has oil worth 36 cents a quart and oil worth 52 cents a quart. How many quarts of each should he mix?
**
1 gal=4 quarts
300 gal=1200 quarts
..
let x=amt of 36ç oil to mix
1200-x=amt of 52ç oil to mix
..
.36x+.52(1200-x)=.4(1200)
.36x+624-.52x=480
.16x=624-480=144
x=144/.16=900
x-900=300
amt of 36ç oil to mix=900 quarts
amt of 52ç oil to mix=300 quarts